(a) Prove that
(−1 + i

3)n + (−1 − i

3)n
has either the value 2n+1 or the value −2
n
if n is any integer (positive, negative or zero).
1
Expert's answer
2020-07-05T17:35:15-0400

(-1+i√3)"^n" + (-1-i√3)"^n"

= 2"^n" ("\frac{-1+i√3}{2})^n" + 2"^n" ("\frac{-1-i√3}{2})^n"

= "2^n" "(\frac{-1}{2}+\frac{i√3}{2})^n" + "2^n" "(\frac{-1}{2}-\frac{i√3}{2})^n"

= "2^n" ( cos "\frac{2π}{3}+i" sin"\frac{2π}{3}" )"^n" + 2"^n" ( cos"\frac{2π}{3}-i"sin"\frac{2π}{3}" )"^n"

= "2^n" ( cos "\frac{2nπ}{3}+" isin"\frac{2nπ}{3}" ) + "2^n" ( cos"\frac{2nπ}{3}-" isin"\frac{2nπ}{3}" ) by D'Moivres theorem

= "2^n" . 2 cos"\frac{2nπ}{3}"

= "2^{n+1}" cos"\frac{2nπ}{3}"

Now as n is an integer, n can be any one of the forms 3k, 3k+1, 3k+2 where k is an integer.

When n = 3k

"cos \frac{2nπ}{3}" = "cos \frac{2(3k)π}{3}" = cos 2kπ = 1

So "(-1+i\u221a3)^n+(-1-i\u221a3)^n"

= 2"^{n+1}"

When n = 3k+1

"cos \frac{2nπ}{3}" = "cos \frac{2(3k+1)π}{3}" = cos(2kπ+"\frac{2π}{3})" = cos"\frac{2π}{3} = cos(π-\frac{π}{3})= -cos\frac{π}{3}" = "-\frac{1}{2}"

So "(-1+i\u221a3)^n+(-1-i\u221a3)^n"

= 2"^{n+1}(-\frac{1}{2})"

= "-2^n"

When n = 3k+2

"cos \frac{2nπ}{3}" = "cos \frac{2(3k+2)π}{3}" = cos(2kπ +"\frac{4π}{3})" = cos"\frac{4π}{3}" = cos(π+"\frac{π}{3}" )= -cos"\frac{π}{3}" = "-\frac{1}{2}"

So "(-1+i\u221a3)^n+(-1-i\u221a3)^n"

= "2^{n+1}(-\frac{1}{2})"

= "-2^n"

Finally we can conclude that

"(-1+i\u221a3)^n+(-1-i\u221a3)^n"

= 2"^{n+1}" or "-" 2"^{n}" if n is an integer (positive, negative or zero)